Crosshill Company’s total overhead costs at various levels of activity are presented below:

Month Machine-Hours Total Overhead Cost
     April 54,000    $188,180  
     May 44,000    $163,980  
     June 64,000    $212,380  
     July 74,000    $236,580  

Assume that the overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 44,000-machine-hour level of activity in May is as follows:

  Utilities (variable) $ 52,800  
  Supervisory salaries (fixed) 44,000
  Maintenance (mixed) 67,180
  Total overhead cost $ 163,980

The company wants to break down the maintenance cost into its variable and fixed cost elements.

1. Estimate how much of the $236,580 of overhead cost in July was maintenance cost. (Hint: To do this, first determine how much of the $236,580 consisted of utilities and supervisory salaries. Think about the behaviour of variable and fixed costs within the relevant range.) (Do not round intermediate calculations.)

2. Using the high–low method, estimate a cost formula for maintenance. (Do not round your intermediate calculations. Round the "Variable cost per unit" to 2 decimal places.)

Machine- Hours Maintenance
Cost High activity level
Low activity level
Change 0 $0
Variable cost element per MH
Fixed cost element

Y= ___+ ____x

3. Express the company’s total overhead cost in the form Y = a + bX. (Do not round your intermediate calculations. Round the "Variable cost per unit" to 2 decimal places.)

Variable Cost per
Machine-Hour Fixed Cost
Utilities cost
Supervisory salaries cost
Maintenance cost
Total overhead cost $0.00 $0

Y= ___+ ____x

4. What total overhead cost would you expect to be incurred at an activity level of 49,000 machine-hours? (Do not round intermediate calculations.)

Total overhead cost =
Answer #1
1
Total overhead cost 236580
Less: Utilities 88800 =52800/44000*74000
Less: Supervisory salaries 44000
Maintenance cost in July 103780
2
Machine hours Maintenance cost
High activity level 74000 103780
Low activity level 44000 67180
Change 30000 36600
Variable cost element 1.22 =36600/30000
Fixed cost element 13500 =103780-(74000*1.22)
Y = 13500 + 1.22 X
3
Variable cost per
machine hour
Fixed cost
Utilities cost 1.20
Supervisory salaries cost 44000
Maintenance cost 1.22 13500
Total overhead cost 2.42 57500
Y = 57500 + 2.42 X
4
Total overhead costs 176080 =57500+(49000*2.42)